A February 1989 issue of Fortune details efforts by companies to increase the rate at which they can develop, produce and market their products. A survey of 50 companies by Kaider Associates, a consulting firm located in Vienna, VA found that almost all firms placed emphasis upon "time-based strategy" (TBS), a strategy found on the idea that "speed kills the competition". One firm in particular, General Education (GE), became concerned about the length of time it took its Akron, OH plant to deliver custom made circuit breaker panels. GE felt that this plant took, on average, about 3 weeks to make such deliveries after receiving an order. To verify this, the Akron, OH plant manager took a sample of 100 recent orders and found the average delivery time to be 3.4 weeks with a standard deviation of 1.1 weeks. At the 2% level, what do these results indicate regarding GE's claim?
Answer #1

z value for the test = x - mean / SD /(n)1/2

= 3.4-3 / 1.1/(100)1/2

= 0.4 /0.11 =3.63

table value for z for 2% level of significance =2.33

Since the calculated value is more than table value of z, the null hypothesis is rejected, and it is evident that there is significant difference between the figures observed and figures claimed.